OpenStudy (anonymous):

Use substitution to find the solutions (if any). x/5 - y/10 = 4 and 3x/4 + 2y/3 = 8

OpenStudy (vijay):

416/25 and -168/25

OpenStudy (anonymous):

Multiply E1 by 10 and E2 by 12. Then you have: 2x - y = 40 and 9x + 8y = 96 This is much easier to solve. Multiply E1 by -9 and E2 by 2. -18x + 9y = -360 18x + 16y = 192 25y = -168 y = -168/25 now you can substitute for x. 2x - (-168/25) = 40 Then x = 416/25 If this helps, please leave a medal. Make sure to check out http://www.tutorsean.net for free math training videos.

OpenStudy (anonymous):

E for effort!